A body travels 15 m from A to B and then moves 20 m at right angle to AB. Draw the motion of the body from initial till final position and Calculate the distance and displacement travelled by the body.
1
Expert's answer
2020-08-02T15:03:54-0400

The total distance is the sum of distances AB and BC, so it is equal to 15 m +20 m = 35 m. But the displacement is equal to the length of vector AC, that can be calculated by the Pythagorean theorem: "AC = \sqrt{AB^2+BC^2} = 25\,\mathrm{m}."